OpenStudy (anonymous):

Factor completely and check mentally 18x^2yz+12xy^2z+6xyz^2

OpenStudy (dan815):

6 is common number.. xyz is common as well

OpenStudy (anonymous):

6 is common because it's a multiple?

OpenStudy (anonymous):

Okay, now what.

OpenStudy (dan815):

okay ill do the middle term this time.. you dothe other 2 terms

OpenStudy (anonymous):

Noooo omg don't do anything. I just need to learn HOW to do it.

OpenStudy (muzzack):

Step by step solution : Simplify \(\large\ 18x^2yz+12xy^2z + (2•3xyz^2)\) Pull out like factors : \(\large\ 18x^2yz + 12xy^2z + 6xyz^2 = 6xyz • (3x + 2y + z)\) Final result : \(\huge\ 6xyz • (3x + 2y + z)\)

OpenStudy (dan815):

6xyz(_____+_______+_______) The middle term ^ has to be 12xy^2z What should you multiply 6xyz by to get 12xy^2z 6xyz * 2 * y = 12xy^2z so... we put 6xyz(_____+__2* y__+_______)

OpenStudy (dan815):

Muzzack explain to him im simpler terms T_T i dont know how

OpenStudy (anonymous):

Wait, how did you figure that out?

OpenStudy (anonymous):

"What should you multiply 6xyz by to get 12xy^2z 6xyz * 2 * y = 12xy^2z so... we put" that part..

OpenStudy (anonymous):

-_-